# DrivingGen: A Comprehensive Benchmark for Generative Video World Models in Autonomous Driving

### v1

@@ 1 Introduction
Driven by scalable learning techniques, generative video models have made remarkable progress in recent years, enabling the synthesis of high-fidelity videos across diverse scenes and motions. These models suggest a promising path toward “world models” – predictive simulators capable of imagining the future, which can support planning, simulation, and decision-making in complex, dynamic environments. Inspired by this vision, there has been an accelerating surge in developing driving world models: generative models specialized for predicting future driving scenarios. Given an initial scene and optional conditions (e.g., text prompts, driving actions), a driving world model predicts both the ego-vehicle’s future movements and the evolution of surrounding agents’ trajectories. Such models enable closed-loop simulation and synthetic data generation, reducing reliance on real-world data and offering a promising means to explore out-of-distribution scenarios safely (gao2024vista; hassan2024gemgeneralizableegovisionmultimodal; mousakhan2025orbisovercomingchallengeslonghorizon; li2025driversenavigationworldmodel; wang2025deployable; zhou2025smartpretrain). Driving world models are also tightly coupled with end-to-end autonomous driving systems, where errors in predicted future scenes and trajectories can directly lead to unsafe decisions (shao2023safety; shao2024lmdrive; shao2023reasonnet; wang2023efficient).
@@
VBench (huang2023vbenchcomprehensivebenchmarksuite)
WorldModelBench (li2025worldmodelbenchjudgingvideogeneration) Instruction
WorldScore (duan2025worldscoreunifiedevaluationbenchmark) Traj.
Vista (gao2024vista) Visual Human eval Traj.
GEM (hassan2024gemgeneralizableegovisionmultimodal) Visual Human eval Human eval, Agent Traj.
Doe-1 (zheng2024doe1closedloopautonomousdriving) Visual VQA, Planning
Drivingdojo (wang2024drivingdojodatasetadvancinginteractive) Visual Traj.
Driverse (li2025driversenavigationworldmodel) Visual Traj.
UniFuture (liang2025seeingfutureperceivingfuture) Visual Perception
VaViM (bartoccioni2025vavimvavamautonomousdriving) Visual Segmentation
GAIA-2 (russell2025gaia2controllablemultiviewgenerative) Visual Visual, Agent
ReSim (yang2025resimreliableworldsimulation) Visual Traj. Planning
ACT-Bench (arai2024actbenchactioncontrollableworld) Instruction, Traj.
@@
Another major limitation in existing benchmarks for driving world models is the lack of diversity along crucial dimensions essential for real-world deployment. 1) First, Weather and Time of Day coverage is heavily skewed: datasets like nuScenes (caesar2020nuscenesmultimodaldatasetautonomous) are dominated by clear-weather, daytime driving, leaving rare but safety-critical conditions (night, snow, fog) underrepresented. 2) Second, Geographic Coverage is limited, often confined to a few cities or countries, which restricts evaluation across varied scene appearance and with local traffic rules. 3) Third, Driving Maneuvers and Interactions rarely capture the full diversity of agent behaviors and complex multi-agent dynamics, such as pedestrians waiting at crosswalks, aggressive driver cut-ins, or dense traffic scenarios (wang2021socially). This lack of diversity makes it difficult to assess whether generative models can handle the wide range of scenarios encountered in real-world driving, undermining their reliability and safety for large-scale deployment.
@@ 3.1 Benchmark Dataset
Generative video models, as a form of world models, offer a promising way to anticipate future driving scenarios, simulate rare or safety-critical events, and ultimately support planning and decision-making. However, real-world driving unfolds under highly variable conditions, encompassing different weather, lighting, regions, and complex maneuvers. Therefore, evaluating generative models across diverse scenarios is crucial to ensure their robustness and reliability. To this end, the majority of existing works (gao2024vista; hassan2024gemgeneralizableegovisionmultimodal; liang2025seeingfutureperceivingfuture; wang2024drivingdojodatasetadvancinginteractive; bartoccioni2025vavimvavamautonomousdriving; wang2024distillnerf) in driving world models mainly utilize nuScenes (caesar2020nuscenesmultimodaldatasetautonomous) and OpenDV (zheng2024genadgenerativeendtoendautonomous) datasets for evaluation. However, the diversity of weather, region, time of day, and driving maneuvers in these datasets is limited and highly biases the data distribution. For example, as shown in Fig. 2(a), over 80% of the nuScenes validation data and 90% of the OpenDV validation data are collected during normal sunny daytime conditions. Additionally, the data are collected from a limited number of vehicles and locations, which further limits the comprehensiveness. Based on this observation, we curated a significantly more diverse dataset. An overview of our dataset is presented in Fig. 2(a) and Fig. 2(b).
@@
  • Ego-Conditioned Track complements the open-domain track. While the open-domain setting evaluates generalization to diverse unseen scenarios, it does not verify whether the generated trajectories follow a specified conditioning trajectory—a property that is critical for robotics and self-driving applications. The ego-conditioned track therefore focuses on trajectory controllability, measuring how well the trajectories derived from generated videos align with the given ego-trajectory instructions. The ego trajectory is optional for model input and only provided in this track. To construct it, we aggregate data from five open-source driving datasets: Zod (zod2023) (Europe), DrivingDojo (wang2024drivingdojodatasetadvancinginteractive) (China), COVLA (covla_wacv2025) (Japan), nuPlan (karnchanachari2024learningbasedplanningthenuplanbenchmark) (US), and WOMD (sun2020scalabilityperceptionautonomousdriving) (US).
Each data sample in the dataset consists of three components: a front-view RGB image (vision), a scene description (language), and an optional ego trajectory (action). For each scene, we employ Qwen (Qwen2.5-VL) to capture descriptions of the future dynamics and camera movements within the scene. Given the time-consuming nature of video generation, we limit the number of samples for efficient testing and iteration, while ensuring quality and diversity. The dataset includes 400 samples—200 per track—striking a balance between efficiency and meaningful evaluation.
@@ 3.2 Benchmark Metrics
For all video models, our DrivingGen metrics cover three key dimensions: distribution, quality, and temporal consistency, evaluated for both videos and trajectories. We extract trajectories using standard PnP method within a SIFT and RANSAC scheme (Lowe2004-pe; Fischler1981-cb; 5995464) and UniDepthV2 (piccinelli2025unidepthv2). We provide the details of our SLAM pipeline (7219438; schoenberger2016sfm; teed2022droidslamdeepvisualslam; qu2024implicit), including guaranteeing that all videos reconstruct trajectories and a discussion to compare other benchmarks’ trajectory reconstruction methods in Appendix B.2. For models conditioned on ego trajectories, we include a fourth dimension: trajectory alignment, measuring adherence to the input. Table 2 lists the metrics, grouped into four categories detailed below, each targeting a different aspect of video fidelity.
@@ 3.2.1 Distribution
How far is the generative distribution from the data distribution? A common practice is to measure Fréchet Video Distance (FVD) (unterthiner2019accurategenerativemodelsvideo) on generated videos. However, our key insight is that video quality is not solely determined by visual realism—equally important, especially for self-driving and embodied agents, is the realism of the induced ego-motion. Focusing only on visual fidelity gives an incomplete picture. Therefore, we evaluate distributional closeness across both videos and trajectories, capturing complementary perspectives from visual perception and robotics.
For the video distribution, we utilize FVD to quantify the similarity between generated videos and real videos. Specifically, we follow the standardized computation protocol from the original StyleGAN-V (skorokhodov2022styleganvcontinuousvideogenerator). For the trajectory distribution, we introduce a novel metric, Fréchet Trajectory Distance (FTD), a distributional metric tailored for evaluating driving trajectories. The key requirement is a trajectory encoder that maps trajectories into a latent space suitable for measuring distributional distance. To this end, we draw from the motion prediction domain—where models themselves are generative of future trajectories, and adopt the encoder of Motion Transformer (MTR) (shi2023motiontransformerglobalintention) as our encoding model. Details of FTD computation are provided in Appendix B.3.
@@ 3.2.2 Quality
Visual Quality. A common practice in generative video evaluation is to assess general perceptual quality with automatic, reference-free estimators aligned with human judgments. Specifically, we adopt CLIP-IQA+ (wang2022exploringclipassessinglook), which leverages CLIP’s vision-language representations to predict perceptual quality scores consistent with human subjective assessments. While effective, such subjective perceptual quality does not always align with what matters for driving, which unfolds outdoors, involves multiple agents, and occurs under real-world constraints. To additionally consider driving-specific imaging quality, we further adopt the Modulation Mitigation Probability (MMP) metric from the IEEE Automotive P2020 standard (ieee2018ieee; 9969530). MMP targets Pulse-Width Modulation (PWM)-induced flicker that can disrupt perception and tracking, and reports the fraction of time windows where residual temporal luminance modulation falls below a small threshold. Implementation details are in Appendix B.4.
@@ 3.2.3 Temporal Consistency
Video Consistency. Existing metrics directly calculate the consistency between consecutive frames (or each frame to the first) at a fixed rate. However, it is easily hackable by generating near-static videos. To measure temporal consistency while accounting for the actual motion in the scene, we first pass the generated videos through an off-the-shelf optical flow model (wang2024searaftsimpleefficientaccurate) to compute the median optical flow magnitude per frame. We then adaptively downsample: videos with lower motion are sampled more sparsely so that the per-step displacement becomes comparable to normal/high-speed driving. After this, the similarity of the DINOv3 (siméoni2025dinov3) features between consecutive frames of the downsampled videos is reported as the video consistency score. Unlike fixed-stride metrics, our approach fairly measures temporal consistency across videos with varying motion speeds, preventing static or near-static videos from obtaining artificially high scores.
Agent Appearance Consistency. Measuring only scene-level features can overlook small temporal changes in individual agents, such as shifts in color, texture, or shape, while these agents are often the key focus for driving, as they would more directly impact driving behavior and safety. To measure the agent’s temporal consistency, we therefore detect agents in the first frame, track them across the video, crop their bounding boxes, and compute consistency purely at the agent level. We use YOLOv10 (wang2024yolov10) as the detector and SAM2 (ravi2024sam2; yang2024samuraiadaptingsegmentmodel) for tracking. We measure DINOv3 feature similarity across consecutive frames and to the first frame.
Agent Abnormal Disappearance. In addition to appearance stability, agents in driving scenes must persist in a physically plausible manner. Sudden, non-physical disappearances of surrounding agents are commonly observed in generated videos, which can compromise realism and safety. DrivingGen quantifies this by diagnosing whether an agent’s disappearance is normal (e.g., leaving the field of view or being occluded) or abnormal. We consider three key frames for each disappearing agent: the first and the last frames where the agent is visible, and the first frame after it vanishes. A vision large language model (VLM) (Qwen2.5-VL; shao2024visual; liu2024sphinx; zong2024mova; li2024llava; qu2025spatialvla; qu2025eo), Cosmos-Reason1 (nvidia2025cosmosreason1physicalcommonsense), is prompted to judge disappearance based on visual and motion continuity, and the agent’s local interactions with surrounding agents. We report the percentage of videos with no abnormal disappearances as the score. Implementation Details can be found in Appendix B.6.
@@ 3.2.4 Trajectory Alignment
Dynamic Time Warping (DTW). In addition to ADE, which compares trajectories at each time step, we introduce a complementary metric that captures the overall contour and shape of the trajectory. Specifically, DTW (Keogh2000ScalingUD) aligns predicted and reference trajectories via non-linear time warping and measures their path-shape discrepancy using Euclidean point-wise cost.
@@ 4 Experiments
Evaluation Setup. We evaluate 14 competitive generative world models on DrivingGen, spanning three categories. 1) First, we include 7 general video world models, comprising two commercial closed-source models, Gen-3 (runway2024gen3) and Kling (kuaishou2024kling), and five well-known open-source models: CogVideoX (yang2024cogvideox), Wan (wan2025wanopenadvancedlargescale), HunyuanVideo (kong2024hunyuanvideo), LTX-Video (HaCohen2024LTXVideo), and SkyReels (chen2025skyreelsv2infinitelengthfilmgenerative). 2) Second, we evaluate 2 physical world models that are developed specifically for the physical robotics domain, Cosmos-Predict1 (agarwal2025cosmos) and Cosmos-Predict2 (cosmos-predict2). 3) Third, we assess 5 driving-specific world models: Vista (gao2024vista), DrivingDojo (wang2024drivingdojodatasetadvancinginteractive), GEM (hassan2024gemgeneralizableegovisionmultimodal), VaViM (bartoccioni2025vavimvavamautonomousdriving), and UniFuture (liang2025seeingfutureperceivingfuture). All models are evaluated on a prediction horizon of 100 frames. We report the time and resource cost for our DrivingGen benchmark in Appendix B.8.
@@ A.1 Generative World Models and their application in Driving
Driven by advances in image generative modeling (kingma2013auto; goodfellow2014generative; esser2021taming; ho2020denoising; peebles2023scalable; zong2024easyref; he2025high), the landscape of large-scale video models has evolved significantly, particularly in diffusion-based frameworks. Closed-source models (videoworldsimulators2024; kuaishou2024kling; luma2024dm; runway2024gen3; pixverse; dreamina; minimax2024hailuo; tongyi2024wanxiang; pika2024pika; shao2024vividface), mainly developed by major technology companies, aim at high-quality, professional video generation with extensive resources invested. Sora (videoworldsimulators2024), introduced by OpenAI, marked a significant leap in Video Generation. Open-source models (rombach2022high; ho2022classifier; hacohen2024ltx; kong2024hunyuanvideo; wan2025wanopenadvancedlargescale; yang2024cogvideox; agarwal2025cosmos), typically based on stable diffusion (rombach2022high) and flow matching (li2025mixgrpounlockingflowbasedgrpo), are quickly expanding and making real contributions to video generation as well. Wan (wan2025wanopenadvancedlargescale), an open-source model, is widely used for video generation and has achieved SOTA results on many benchmarks. Recent years have also seen remarkable progress in both multimodal understanding and generation models (li2025unif2aceunifiedfinegrainedface; zhang2025unifiedmultimodalunderstandinggeneration).
Besides general video generation, driving-focused generative models use sensor data such as lidar point clouds (zheng2024occworld; yang2024visual) or images (gao2024vista; hassan2024gemgeneralizableegovisionmultimodal; hu2023gaia; wang2024drivedreamer; wang2024driving; yang2024generalized; zhao2025drivedreamer). Since this work emphasizes video generation, we focus on image-based methods. Early approaches before Vista (gao2024vista) rely on multi-view RGB inputs and high-definition maps or 3D boxes, limiting generalization to new datasets and open-domain videos. Vista-based methods (hassan2024gemgeneralizableegovisionmultimodal; li2025driversenavigationworldmodel; mousakhan2025orbisovercomingchallengeslonghorizon) simplify inputs to a single front-view image with optional ego trajectories, improving scalability to YouTube videos and enabling broader open-domain evaluation.

### v2

@@ 1 Introduction
Driven by scalable learning techniques, generative video models have made remarkable progress in recent years, enabling the synthesis of high-fidelity videos across diverse scenes and motions. These models suggest a promising path toward “world models” – predictive simulators capable of imagining the future, which can support planning, simulation, and decision-making in complex, dynamic environments. Inspired by this vision, there has been an accelerating surge in developing driving world models: generative models specialized for predicting future driving scenarios. Given an initial scene and optional conditions (e.g., text prompts, driving actions), a driving world model predicts both the ego-vehicle’s future movements and the evolution of surrounding agents’ trajectories. Such models enable closed-loop simulation and synthetic data generation, reducing reliance on real-world data and offering a promising means to explore out-of-distribution scenarios safely (Gao et al., 2024; Hassan et al., 2024; Mousakhan et al., 2025; Li et al., 2025d; Wang et al., 2025; Zhou et al., 2025). Driving world models are also tightly coupled with end-to-end autonomous driving systems, where errors in predicted future scenes and trajectories can directly lead to unsafe decisions (Shao et al., 2023a; 2024a; b; Wang et al., 2023).
@@
VBench (Huang et al., 2023)
WorldModelBench (Li et al., 2025a) Instruction
WorldScore (Duan et al., 2025) Traj.
Vista (Gao et al., 2024) Visual Human eval Traj.
GEM (Hassan et al., 2024) Visual Human eval Human eval, Agent Traj.
Doe-1 (Zheng et al., 2024c) Visual VQA, Planning
Drivingdojo (Wang et al., 2024e) Visual Traj.
Driverse (Li et al., 2025d) Visual Traj.
UniFuture (Liang et al., 2025) Visual Perception
VaViM (Bartoccioni et al., 2025) Visual Segmentation
GAIA-2 (Russell et al., 2025) Visual Visual, Agent
ReSim (Yang et al., 2025) Visual Traj. Planning
ACT-Bench (Arai et al., 2024) Instruction, Traj.
@@
Another major limitation in existing benchmarks for driving world models is the lack of diversity along crucial dimensions essential for real-world deployment. 1) First, Weather and Time of Day coverage is heavily skewed: datasets like nuScenes (Caesar et al., 2020) are dominated by clear-weather, daytime driving, leaving rare but safety-critical conditions (night, snow, fog) underrepresented. 2) Second, Geographic Coverage is limited, often confined to a few cities or countries, which restricts evaluation across varied scene appearance and with local traffic rules. 3) Third, Driving Maneuvers and Interactions rarely capture the full diversity of agent behaviors and complex multi-agent dynamics, such as pedestrians waiting at crosswalks, aggressive driver cut-ins, or dense traffic scenarios (Wang et al., 2021). This lack of diversity makes it difficult to assess whether generative models can handle the wide range of scenarios encountered in real-world driving, undermining their reliability and safety for large-scale deployment.
@@ 3.1 Benchmark Dataset
Generative video models, as a form of world models, offer a promising way to anticipate future driving scenarios, simulate rare or safety-critical events, and ultimately support planning and decision-making. However, real-world driving unfolds under highly variable conditions, encompassing different weather, lighting, regions, and complex maneuvers. Therefore, evaluating generative models across diverse scenarios is crucial to ensure their robustness and reliability. To this end, the majority of existing works (Gao et al., 2024; Hassan et al., 2024; Liang et al., 2025; Wang et al., 2024e; Bartoccioni et al., 2025; Wang et al., 2024b) in driving world models mainly utilize nuScenes (Caesar et al., 2020) and OpenDV (Zheng et al., 2024b) datasets for evaluation. However, the diversity of weather, region, time of day, and driving maneuvers in these datasets is limited and highly biases the data distribution. For example, as shown in Fig. 2(a), over 80% of the nuScenes validation data and 90% of the OpenDV validation data are collected during normal sunny daytime conditions. Additionally, the data are collected from a limited number of vehicles and locations, which further limits the comprehensiveness. Based on this observation, we curated a significantly more diverse dataset. An overview of our dataset is presented in Fig. 2(a) and Fig. 2(b).
@@
  • Ego-Conditioned Track complements the open-domain track. While the open-domain setting evaluates generalization to diverse unseen scenarios, it does not verify whether the generated trajectories follow a specified conditioning trajectory—a property that is critical for robotics and self-driving applications. The ego-conditioned track therefore focuses on trajectory controllability, measuring how well the trajectories derived from generated videos align with the given ego-trajectory instructions. The ego trajectory is optional for model input and only provided in this track. To construct it, we aggregate data from five open-source driving datasets: Zod (Alibeigi et al., 2023) (Europe), DrivingDojo (Wang et al., 2024e) (China), COVLA (Arai et al., 2025) (Japan), nuPlan (Karnchanachari et al., 2024) (US), and WOMD (Sun et al., 2020) (US).
Each data sample in the dataset consists of three components: a front-view RGB image (vision), a scene description (language), and an optional ego trajectory (action). For each scene, we employ Qwen (Bai et al., 2025) to capture descriptions of the future dynamics and camera movements within the scene. Given the time-consuming nature of video generation, we limit the number of samples for efficient testing and iteration, while ensuring quality and diversity. The dataset includes 400 samples—200 per track—striking a balance between efficiency and meaningful evaluation.
@@ 3.2 Benchmark Metrics
For all video models, our DrivingGen metrics cover three key dimensions: distribution, quality, and temporal consistency, evaluated for both videos and trajectories. We extract trajectories using standard PnP method within a SIFT and RANSAC scheme (Lowe, 2004; Fischler & Bolles, 1981; Kneip et al., 2011) and UniDepthV2 (Piccinelli et al., 2025). We provide the details of our SLAM pipeline (Mur-Artal et al., 2015; Schönberger & Frahm, 2016; Teed & Deng, 2022; Qu et al., 2024), including guaranteeing that all videos reconstruct trajectories and a discussion to compare other benchmarks’ trajectory reconstruction methods in Appendix B.2. For models conditioned on ego trajectories, we include a fourth dimension: trajectory alignment, measuring adherence to the input. Table 2 lists the metrics, grouped into four categories detailed below, each targeting a different aspect of video fidelity.
@@ 3.2.1 Distribution
How far is the generative distribution from the data distribution? A common practice is to measure Fréchet Video Distance (FVD) (Unterthiner et al., 2019) on generated videos. However, our key insight is that video quality is not solely determined by visual realism—equally important, especially for self-driving and embodied agents, is the realism of the induced ego-motion. Focusing only on visual fidelity gives an incomplete picture. Therefore, we evaluate distributional closeness across both videos and trajectories, capturing complementary perspectives from visual perception and robotics.
For the video distribution, we utilize FVD to quantify the similarity between generated videos and real videos. Specifically, we follow the standardized computation protocol from the original StyleGAN-V (Skorokhodov et al., 2022). For the trajectory distribution, we introduce a novel metric, Fréchet Trajectory Distance (FTD), a distributional metric tailored for evaluating driving trajectories. The key requirement is a trajectory encoder that maps trajectories into a latent space suitable for measuring distributional distance. To this end, we draw from the motion prediction domain—where models themselves are generative of future trajectories, and adopt the encoder of Motion Transformer (MTR) (Shi et al., 2023) as our encoding model. Details of FTD computation are provided in Appendix B.3.
@@ 3.2.2 Quality
Visual Quality. A common practice in generative video evaluation is to assess general perceptual quality with automatic, reference-free estimators aligned with human judgments. Specifically, we adopt CLIP-IQA+ (Wang et al., 2022), which leverages CLIP’s vision-language representations to predict perceptual quality scores consistent with human subjective assessments. While effective, such subjective perceptual quality does not always align with what matters for driving, which unfolds outdoors, involves multiple agents, and occurs under real-world constraints. To additionally consider driving-specific imaging quality, we further adopt the Modulation Mitigation Probability (MMP) metric from the IEEE Automotive P2020 standard (IEEE P2020 Working Group, 2018; IEEE, 2022). MMP targets Pulse-Width Modulation (PWM)-induced flicker that can disrupt perception and tracking, and reports the fraction of time windows where residual temporal luminance modulation falls below a small threshold. Implementation details are in Appendix B.4.
@@ 3.2.3 Temporal Consistency
Video Consistency. Existing metrics directly calculate the consistency between consecutive frames (or each frame to the first) at a fixed rate. However, it is easily hackable by generating near-static videos. To measure temporal consistency while accounting for the actual motion in the scene, we first pass the generated videos through an off-the-shelf optical flow model (Wang et al., 2024d) to compute the median optical flow magnitude per frame. We then adaptively downsample: videos with lower motion are sampled more sparsely so that the per-step displacement becomes comparable to normal/high-speed driving. After this, the similarity of the DINOv3 (Siméoni et al., 2025) features between consecutive frames of the downsampled videos is reported as the video consistency score. Unlike fixed-stride metrics, our approach fairly measures temporal consistency across videos with varying motion speeds, preventing static or near-static videos from obtaining artificially high scores.
Agent Appearance Consistency. Measuring only scene-level features can overlook small temporal changes in individual agents, such as shifts in color, texture, or shape, while these agents are often the key focus for driving, as they would more directly impact driving behavior and safety. To measure the agent’s temporal consistency, we therefore detect agents in the first frame, track them across the video, crop their bounding boxes, and compute consistency purely at the agent level. We use YOLOv10 (Wang et al., 2024a) as the detector and SAM2 (Ravi et al., 2024; Yang et al., 2024a) for tracking. We measure DINOv3 feature similarity across consecutive frames and to the first frame.
Agent Abnormal Disappearance. In addition to appearance stability, agents in driving scenes must persist in a physically plausible manner. Sudden, non-physical disappearances of surrounding agents are commonly observed in generated videos, which can compromise realism and safety. DrivingGen quantifies this by diagnosing whether an agent’s disappearance is normal (e.g., leaving the field of view or being occluded) or abnormal. We consider three key frames for each disappearing agent: the first and the last frames where the agent is visible, and the first frame after it vanishes. A vision large language model (VLM) (Bai et al., 2025; Shao et al., 2024b; Liu et al., 2024; Zong et al., 2024b; Li et al., 2024; Qu et al., 2025b; a), Cosmos-Reason1 (NVIDIA et al., 2025), is prompted to judge disappearance based on visual and motion continuity, and the agent’s local interactions with surrounding agents. We report the percentage of videos with no abnormal disappearances as the score. Implementation Details can be found in Appendix B.6.
@@ 3.2.4 Trajectory Alignment
Dynamic Time Warping (DTW). In addition to ADE, which compares trajectories at each time step, we introduce a complementary metric that captures the overall contour and shape of the trajectory. Specifically, DTW (Keogh & Pazzani, 2000) aligns predicted and reference trajectories via non-linear time warping and measures their path-shape discrepancy using Euclidean point-wise cost.
@@ 4 Experiments
Evaluation Setup. We evaluate 14 competitive generative world models on DrivingGen, spanning three categories. 1) First, we include 7 general video world models, comprising two commercial closed-source models, Gen-3 (Runway, 2024) and Kling (Kuaishou, 2024), and five well-known open-source models: CogVideoX (Yang et al., 2024e), Wan (Wan et al., 2025), HunyuanVideo (Kong et al., 2024), LTX-Video (HaCohen et al., 2024), and SkyReels (Chen et al., 2025). 2) Second, we evaluate 2 physical world models that are developed specifically for the physical robotics domain, Cosmos-Predict1 (Agarwal et al., 2025) and Cosmos-Predict2 (NVIDIA Cosmos, 2025). 3) Third, we assess 5 driving-specific world models: Vista (Gao et al., 2024), DrivingDojo (Wang et al., 2024e), GEM (Hassan et al., 2024), VaViM (Bartoccioni et al., 2025), and UniFuture (Liang et al., 2025). All models are evaluated on a prediction horizon of 100 frames. We report the time and resource cost for our DrivingGen benchmark in Appendix B.8.
@@ A.1 Generative World Models and their application in Driving
Driven by advances in image generative modeling (Kingma & Welling, 2013; Goodfellow et al., 2014; Esser et al., 2021; Ho et al., 2020; Peebles & Xie, 2023; Zong et al., 2024a; He et al., 2025), the landscape of large-scale video models has evolved significantly, particularly in diffusion-based frameworks. Closed-source models (Brooks et al., 2024; Kuaishou, 2024; Luma Labs, 2024; Runway, 2024; PixVerse, 2023; CapCut, 2024; MiniMax, 2024; Pika Labs, 2024; Shao et al., 2024c), mainly developed by major technology companies, aim at high-quality, professional video generation with extensive resources invested. Sora (Brooks et al., 2024), introduced by OpenAI, marked a significant leap in Video Generation. Open-source models (Rombach et al., 2022; Ho & Salimans, 2022; HaCohen et al., 2024; Kong et al., 2024; Wan et al., 2025; Yang et al., 2024e; Agarwal et al., 2025), typically based on stable diffusion (Rombach et al., 2022) and flow matching (Li et al., 2025b), are quickly expanding and making real contributions to video generation as well. Wan (Wan et al., 2025), an open-source model, is widely used for video generation and has achieved SOTA results on many benchmarks. Recent years have also seen remarkable progress in both multimodal understanding and generation models (Li et al., 2025c; Zhang et al., 2025).
Besides general video generation, driving-focused generative models use sensor data such as lidar point clouds (Zheng et al., 2024a; Yang et al., 2024d) or images (Gao et al., 2024; Hassan et al., 2024; Hu et al., 2023; Wang et al., 2024c; f; Yang et al., 2024b; Zhao et al., 2024). Since this work emphasizes video generation, we focus on image-based methods. Early approaches before Vista (Gao et al., 2024) rely on multi-view RGB inputs and high-definition maps or 3D boxes, limiting generalization to new datasets and open-domain videos. Vista-based methods (Hassan et al., 2024; Li et al., 2025d; Mousakhan et al., 2025) simplify inputs to a single front-view image with optional ego trajectories, improving scalability to YouTube videos and enabling broader open-domain evaluation.